2ydx+x(x^2lny-1)dy=0

1
Expert's answer
2022-03-25T06:09:34-0400

"\frac{2y}{x^3}dx+\left( \ln y-\frac{1}{x^2} \right) dy=0\\d\left( -\frac{y}{x^2} \right) +\ln ydy=0\\d\left( -\frac{y}{x^2}+y\ln y-y \right) =0\\-\frac{y}{x^2}+y\ln y-y=Const"